OpenStudy (anonymous):

what is the distance between P(0,0) and Q(5,12)

OpenStudy (aaliyahfairgood):

PQ=root [(0-5)^2 +(0-12)^2] PQ=root [(5)^2+(12)^2] PQ=root[25+144] PQ=root(21361) PQ=146.1540283

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy @hayleymarie4280 ! :) \[distance formula: \sqrt{(x _{2}-x _{1})^{2} + (y _{2}-y _{1})^{2}}\]when given points (x1, y1) and (x2, y2) so plug in x2 = 5 x1 = 0 y2 = 12 y1 = 0 and you will get \[\sqrt{(5-0)^2+(12-0)^2}~=~~?\]can you solve from there? ^_^